You pull two cards, one at a time, from a deck of cards, without replacement. What is the probability that the second card you pick has a different color, or different suit, than the first card
Answer by greenestamps(13327)



When the second card is drawn, there are 51 cards left in the deck.

The opposite of the desired outcome is that the second card is the same color and the same suit as the first -- i.e., it is the same suit as the first. There are 12 cards of the same suit remaining, so the probability of NOT getting the desired outcome is 12/51 = 4/17.

So the probability of getting the desired outcome is 1 - 4/17 = 13/17.

ANSWER: 13/17